OpenStudy (anonymous):

Minimize S = x + 19y with xy = 19 and both x and y > 0.

OpenStudy (anonymous):

S = 19/y + 19y

OpenStudy (anonymous):

differentiate S w.r.t y and set it equal to 0

OpenStudy (anonymous):

S' = -19/y^2 + 19 = 0

OpenStudy (anonymous):

y^2 = 1

OpenStudy (anonymous):

y = 1 or -1

OpenStudy (anonymous):

y can't be negative, so y = 1

OpenStudy (anonymous):

x = 19